OpenStudy (anonymous):

What are the two cases in which the Laws of Sines can be applied to solve a non-right triangle? Case I. You know the measures of two angles and any side of the triangle. Case II. You know the measures of two sides and an angle opposite one of the two known sides. Case III. You know the measures of two sides and the included angle. Case IV. You know the measures of all three sides. A. Case I and Case III B. Case II and Case IV C. Case III and Case IV D. Case I and Case II

OpenStudy (anonymous):

use a/sinA=b/sinB=c/sinC.....here a,b,c are sides and A,B,C are angles

OpenStudy (anonymous):

i hate this question, it is a dumb thing to memorize. if you have a triangle and you want to solve it, the easiest way is to try and use the law of sines, \[\frac{\sin(A)}{a}=\frac{\sin(B)}{b}=\frac{\sin(C)}{c}\] but in order to do this you need two out of the three numbers if you have them, you can do it, if you don't then you cannot

OpenStudy (anonymous):

is it B

OpenStudy (anonymous):

it is not B

OpenStudy (anonymous):

or D

OpenStudy (anonymous):

if you know only the three sides, then you do not know two out of the three numbers

OpenStudy (anonymous):

its c

OpenStudy (anonymous):

yes, it is D and you should draw a picture to see that if you are in case I then you know three out of the 4 numbers here, likewise for case II \[\frac{\sin(A)}{a}=\frac{\sin(B)}{b}\]

OpenStudy (anonymous):

and the reason it is true for case I is that if you know two angles, you also know the third, since they have to add up to 180

OpenStudy (waleed_imtiaz):

I think its A...... For case || and |V we can't use law of sines........

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

yes you can. draw the picture and you will see it

OpenStudy (waleed_imtiaz):

okay let me think for a moment again.....

OpenStudy (anonymous):

this is the law of sines \[\frac{\sin(A)}{a}=\frac{\sin(B)}{b}\]

OpenStudy (anonymous):

in order for this to be of use, you have to know three out of the four numbers

OpenStudy (waleed_imtiaz):

Its not D I think so...... Because @satellite73 if we know two sides and the angle opposite to both the sides ....... Then we can't solve it..

OpenStudy (anonymous):

if you know the angle opposite one of the sides, then if you call the side \(a\) and the angle opposite that side \(A\) then you know \[\frac{\sin(A)}{a}\]

OpenStudy (anonymous):

and if you know some other side, call it \(b\) then you know three out of the four numbers in \[\frac{\sin(A)}{a}=\frac{\sin(B)}{b}\] so you can find the fourth

OpenStudy (anonymous):

sorry messed it up a bit, yeah it's d

OpenStudy (anonymous):

but as i said before, this is really a bad way to think about the law of sines. you should use it if you can, and you can tell easily if you can use it, by checking to see if you have three out of the four numbers

OpenStudy (anonymous):

so what if i was law of cosine then it would be C

OpenStudy (anonymous):

youhave six variables you know a,b,c i.e sides from case IV and two out of three angles A,B,C and then only one unknown angle remains which could be found by the sine law i had mentioned earlier.....hope it makes the picture clear

OpenStudy (waleed_imtiaz):

oh,okay..... I got it.thanks @satellite73 ..... :)

OpenStudy (anonymous):

@akash809 if you know two out of three angles you would use the fact they add to 180

OpenStudy (anonymous):

yeah noticed that there is no option of case I and Case IV and wouldn't make much sense too ,looks like i m not reading the options properly :)...time to sleep now